OpenStudy (anonymous):

Solve and Graph -10x=5y

OpenStudy (anonymous):

Do you know how ot find the X and Y intercepts?

OpenStudy (anonymous):

Say you want to find the X intercept, lets set Y equal to 0. -10x=5y -10x=5(0) -10x=0 /-10 x=0 Coordinate 1= (0, 0) Now lets find the Y intercept, lets find X equal to 0. -10x=5y -10(0)=5y 0=5y /5 y=undefined.

OpenStudy (anonymous):

Thanks Goodman. But how will the graph look like.

OpenStudy (anonymous):



OpenStudy (anonymous):

THANKS GOODMAN